# X2C: A Dataset Featuring Nuanced Facial Expressions for Realistic Humanoid Imitation

### v1

@@ 1 Introduction
     Dataset Dataset Characteristics Dataset Quality
Size Asymmetric Expressions Input Dimensionality Annotation Dimensionality Annotation Accuracy Data Alignment Data Diversity
X2C 100,000 $512\times 512\times 3$ 30 ★★★★★ ★★★★★
Smile chen2021smile 15,000 $480\times 320\times 3$ 10 ★★★★ ★★★
Coexpression hu2024human 1000 $113\times 2$ 11 ★★★★ ★★★
@@ 2.1 The Humanoid Robot and Control Values
Jaw Pitch Jaw Yaw Lip Bottom Curl Lip Bottom Depress Left Lip Bottom Depress Middle Lip Bottom Depress Right $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ Lip Corner Raise Left Lip Corner Raise Right Lip Corner Stretch Left Lip Corner Stretch Right Lip Top Curl Lip Top Raise Left $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ Lip Top Raise Middle Lip Top Raise Right Nose Wrinkle Brow Inner Left Brow Inner Right Brow Outer Left $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ Brow Outer Right Eyelid Lower Left Eyelid Lower Right Eyelid Upper Left Eyelid Upper Right Gaze Target Phi $[0,1]$ $[-1,2]$ $[-1,2]$ $[-1,2]$ $[-1,2]$ $[-2.3,2.3]$ Gaze Target Theta Head Pitch Head Roll Head Yaw Neck Pitch Neck Roll $[-1.1,1.1]$ $[-0.5,0.3]$ $[-0.3,0.3]$ $[-0.5,0.5]$ $[-0.3,0.5]$ $[-0.3,0.3]$
@@ 2.3 Sampling and Annotation
We filmed videos of all curated animations using the same device (MacBook Air, 2020) and under consistent configurations (i.e., frame rate, and upper-body pose of the robot) to ensure data consistency. Images were then sampled at a constant timestep of $s=0.02$ seconds and resized to a uniform resolution of 512$\times$512 pixels. To ensure data uniqueness, the resulting images were processed with a structural similarity check to identify and remove near-duplicate frames if their similarity exceeded a threshold ($\theta=0.99$). Given two image patches $x$ and $y$, the Structural Similarity Index (SSIM) is computed as follows wang2004image :
@@ 2.4 Dataset Overview
JP JY LBC LBDL LBDM LBDR LCRL LCRR LCSL LCSR LTC LTRL LTRM LTRR NW $\mu$ 0.635 0.519 0.544 0.565 0.465 0.560 0.355 0.542 0.646 0.403 0.543 0.649 0.540 0.620 0.191 $\sigma$ 0.366 0.190 0.123 0.085 0.087 0.064 0.129 0.087 0.119 0.095 0.153 0.134 0.182 0.140 0.287 $V_{\text{max}}$ 1.000 1.000 0.991 0.750 0.759 0.822 0.802 0.976 1.000 0.991 1.000 1.000 1.000 1.000 1.000 $V_{\text{min}}$ 0.000 0.000 0.000 0.148 0.000 0.185 0.000 0.250 0.295 0.300 0.000 0.475 0.300 0.409 0.000 ${V}_{\text{neu}}$ 1.000 0.500 0.460 0.560 0.430 0.540 0.470 0.620 0.640 0.310 0.410 0.480 0.300 0.450 0.000 BIL BIR BOL BOR ELL ELR EUL EUR GTP GTT HP HR HY NP NR $\mu$ 0.605 0.655 0.613 0.598 1.110 1.060 0.989 0.978 0.074 0.045 0.004 0.005 -0.002 0.008 0.002 $\sigma$ 0.236 0.233 0.211 0.216 0.521 0.518 0.311 0.331 0.311 0.113 0.036 0.021 0.048 0.024 0.009 $V_{\text{max}}$ 1.000 1.000 1.000 1.000 2.000 2.000 2.000 2.000 2.269 1.082 0.328 0.300 0.355 0.204 0.084 $V_{\text{min}}$ 0.000 0.000 0.000 $0$1.000 -1.000 -1.000 -1.000 -2.269 -0.826 -0.371 -0.173 -0.413 -0.104 -0.158 ${V}_{\text{neu}}$ 0.500 0.500 0.500 0.500 1.000 1.000 1.000 1.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000

### v2

@@ 1 Introduction
     Dataset
Dataset Characteristics Dataset Quality
Size
Asymmetric
Expressions
Input
Dimensionality
Annotation
Dimensionality
Annotation
Accuracy
Data
Alignment
Data
Diversity
X2C
100,000 $512\times 512\times 3$ 30 ★★★★★ ★★★★★
Smile chen2021smile
15,000 $480\times 320\times 3$ 10 ★★★★ ★★★
Coexpression hu2024human
1000 $13\times 2 11$ ★★★★ ★★★
@@ 2.1 The Humanoid Robot and Control Values
Jaw Pitch Jaw Yaw Lip Bottom Curl Lip Bottom Depress Left Lip Bottom Depress Middle Lip Bottom Depress Right $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ Lip Corner Raise Left Lip Corner Raise Right Lip Corner Stretch Left Lip Corner Stretch Right Lip Top Curl Lip Top Raise Left $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ Lip Top Raise Middle Lip Top Raise Right Nose Wrinkle Brow Inner Left Brow Inner Right Brow Outer Left $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ $[0,1]$ Brow Outer Right Eyelid Lower Left Eyelid Lower Right Eyelid Upper Left Eyelid Upper Right Gaze Target Phi $[0,1]$ [1,2][-1,2] [1,2][-1,2] [1,2][-1,2] [1,2][-1,2] [2.3,2.3][-2.3,2.3] Gaze Target Theta Head Pitch Head Roll Head Yaw Neck Pitch Neck Roll [1.1,1.1][-1.1,1.1] [0.5,0.3][-0.5,0.3] [0.3,0.3][-0.3,0.3] [0.5,0.5][-0.5,0.5] [0.3,0.5][-0.3,0.5] [0.3,0.3][-0.3,0.3]
@@ 2.3 Sampling and Annotation
We filmed videos of all curated animations using the same device (MacBook Air, 2020) and under consistent configurations (i.e., frame rate, and upper-body pose of the robot) to ensure data consistency. Images were then sampled at a constant timestep of $s=0.05$ seconds and resized to a uniform resolution of $\times512$ pixels. To ensure data uniqueness, the resulting images were processed with a structural similarity check to identify and remove near-duplicate frames if their similarity exceeded a threshold ($\theta=0.99$). Given two image patches $x$ and $y$, the Structural Similarity Index (SSIM) is computed as follows wang2004image :
@@ 2.4 Dataset Overview
JP JY LBC LBDL LBDM LBDR LCRL LCRR LCSL LCSR LTC LTRL LTRM LTRR NW $\mu$ 0.635 0.519 0.544 0.565 0.465 0.560 0.355 0.542 0.646 0.403 0.543 0.649 0.540 0.620 0.191 $\sigma$ 0.366 0.190 0.123 0.085 0.087 0.064 0.129 0.087 0.119 0.095 0.153 0.134 0.182 0.140 0.287 $V_{\text{max}}$ 1.000 1.000 0.991 0.750 0.759 0.822 0.802 0.976 1.000 0.991 1.000 1.000 1.000 1.000 1.000 $V_{\text{min}}$ 0.000 0.000 0.000 0.148 0.000 0.185 0.000 0.250 0.295 0.300 0.000 0.475 0.300 0.409 0.000 ${V}_{\text{neu}}$ 1.000 0.500 0.460 0.560 0.430 0.540 0.470 0.620 0.640 0.310 0.410 0.480 0.300 0.450 0.000 BIL BIR BOL BOR ELL ELR EUL EUR GTP GTT HP HR HY NP NR $\mu$ 0.605 0.655 0.613 0.598 1.110 1.060 0.989 0.978 0.074 0.045 0.004 0.005 -0.002 0.008 0.002 $\sigma$ 0.236 0.233 0.211 0.216 0.521 0.518 0.311 0.331 0.311 0.113 0.036 0.021 0.048 0.024 0.009 $V_{\text{max}}$ 1.000 1.000 1.000 1.000 2.000 2.000 2.000 2.000 2.269 1.082 0.328 0.300 0.355 0.204 0.084 $V_{\text{min}}$ 0.000 0.000 0.000 0.000 -1.000 -1.000 -1.000 -1.000 -2.269 -0.826 -0.371 -0.173 -0.413 -0.104 -0.158 ${V}_{\text{neu}}$ 0.500 0.500 0.500 0.500 1.000 1.000 1.000 1.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000